log base 1/5 ((x^2)+x) - log base 1/5 ((x^2)-x) = -2
x=?

Answer by stanbon(75887)

log base 1/5 ((x^2)+x) - log base 1/5 ((x^2)-x) = -2
------------------
log1/5[(x^2+x)/(x^2-x)] = -2
------------------
(x^2+x)/(x^2-x) = (1/5)^-2
-----
(x+1)/(x-1) = 25
-----
x+1 = 25x-25
24x = 26
x = 13/12
================
Cheers,
Stan H.
================